# Renormalization of the Quantum Stress Tensor Fluctuations and the Limits of Semiclassical Gravity

### v1

@@ Abstract
We analyze the expectation value of the energy–momentum tensor and its fluctuations in quantum field theory on curved spacetimes $\langle T_{ab}\rangle$. A necessary condition for the conceptual consistency of semiclassical gravity, where $\langle T_{ab}\rangle$ represent the sources of the Einstein equations, is that the fluctuations of the energy momentum tensor remain small compared to its expectation value. We study the renormalization of both the energy–momentum tensor $\langle T_{ab}(x)\rangle_{\rm ren}$ and the fluctuation tensor $\langle T_{ab}(x)T_{cd}(x)\rangle_{\rm ren}$ for suitable Hadamard states, using the operator product expansion for a free scalar field on a fixed curved background. We show that squeezed vacua—arising naturally in black hole evaporation and in inflationary cosmology—fail to satisfy the semiclassicality criterion.
@@
In the absence of a consistent and well-defined theory of quantum gravity, semiclassical gravity—defined by Einstein’s equations coupled to the expectation value of the energy–momentum tensor of matter in a quantum state—is expected to provide useful insights into phenomena at the interface between gravity and quantum physics. Even though it is not entirely clear whether the dynamical equations of semiclassical gravity can be given a fully consistent interpretation—owing to questions about well-posedness in generic situations stemming from their higher-derivative character Flanagan:1996gw ; Juarez-Aubry:2022qdp —the semiclassical framework nonetheless plays a central role in conceptual discussions. These range from the physics of black-hole evaporation via Hawking radiation, to analyses of the generalized second law, the fate of information, and various issues related to the emergence of structure in cosmology.
In order to explore the regime of applicability of semiclassical gravity, it is necessary to evaluate the quantum uncertainties of the energy–momentum tensor in comparison to the expectation value of that same quantity. The issue would be straightforward were it not for the fact that both the expectation value and the fluctuations are divergent notions that require suitable renormalization in order to yield finite and meaningful results.
The renormalization of expectation values of operators that are quadratic in the fundamental fields and their derivatives (such as the energy–momentum tensor of free fields) is customarily based on the Hadamard subtraction. An axiomatic analysis proposed by Wald Wald:1995yp assures that the recipe is unique up to the higher order curvature terms which (aside from a finite number of them) are naturally expected to be suppressed by appropriate powers of the Planck scale.
@@
While our results agree in general lines with these earlier works and in the specific settings they considered, our approach offers a systematic and fully covariant framework that is, in principle, applicable to general spacetimes and general (Hadamard) states. Moreover, the OPE perspective ensures automatically that flat spacetime with matter fields in the Minkowski vacuum state is a consistent solution of semiclassical gravity (see (19)).
@@
${{\mathfs{T}}}_{ab^{\prime}}$ $=$ $\nabla_{a}\nabla_{b^{\prime}}-\frac{1}{2}g_{ab}g^{cd^{\prime}}\nabla_{c}\nabla_{d^{\prime}}-\frac{1}{2}g_{ab}m^{2}$ (3)
where $g_{a}^{b^{\prime}}$ represents the parallel transport operators along the unique geodesic connecting the points $x$ and $x^{\prime}$ within a common convex neighbourhood, and $H(x,x^{\prime})$—which in equation (2) corresponds to $C(x,x^{\prime})_{\rm uv}$—is the Hadamard bi-distributional kernel, constructed from geometrically defined quantities, which is state-independent and captures the UV behavior of the (Hadamard) states considered physically meaningful.
@@
In defining the fluctuations of the energy momentum tensor in a given quantum state of the scalar field, the strategy will be to start with the 4-point function $\langle\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi\rangle$, from which we can define multi-tensorial quantity
@@
The OPE is very general and allows for the renormalization of the expectation value of arbitrary products of fields and their derivatives at a same point. However even when the framework ensures the success of the renormalization program, the specific calculations of quantities of interest, it can become difficult in concrete practical applications.
@@
As shown in the appendix A, the 4-point fucntion of the scalar field can be written as
@@
where
and the finite renormalized component of the 4-point function is given by
$\!\!\!\!\!\!\!\!\!\braket{0|\phi(x_{1})\phi(x_{2})|0}_{\rm ren}\braket{0|\phi(x_{3})\phi(x_{4})|0}_{\rm ren}+{\rm other\ pairings}.{}$
Performing the multiple Taylor expansion of the renormalized quantities (those with the subindex “ren”) in (6) (for example writing $\braket{\psi|\phi(x)\phi(x+y)|\psi}_{\rm ren}=\braket{\psi|\phi(x)^{2}|\psi}_{\rm ren}+y^{\mu}\partial_{\mu}\braket{\psi|\phi(x)^{2}|\psi}_{\rm ren}+\cdots$) we would obtain the OPE for the 4-point function of a free field on a curved background. As each of the divergent terms in (6) are canonical, the finite term (8) does not depend on the choice of the vacuum state used in the normal ordering. The renormalized square of the energy momentum tensor can now be given a meaning as
@@
Focusing on the trace, (8) implies that for any squeezed state (defined as a Bogoliubov transformation of an arbitrary vacuum state) such that $\braket{\phi^{2}}_{\rm ren}\not=0$, we have
@@ I Explicit calculation in 2d
$\braket{\psi|T_{\pm\pm}(x)|\psi}_{\rm ren}=\braket{\psi|:T_{\pm\pm}:_{x}|\psi}+\frac{\hbar}{12\pi}\tau_{\pm}(\Omega),$ (11)
@@
$\tau_{\pm}(\Omega)\equiv\left(\partial^{2}_{\pm}\log(\Omega)-\partial_{\pm}\log(\Omega)\partial_{\pm}\log(\Omega)\right),$ (12)
@@
where, as in (I), we denote by $_{X}$ the normal ordered prescription based on the vacuum notion constructed out of the positive frequency notion associated with freely falling coordinates null coordinates $X^{\pm}$. In Appendix B we explicitly show that the previous procedure is equivalent to the OPE perspective advocated in Hollands:2023txn . We also show that the relevant renormalized fluctuations are given by
@@
$\frac{\hbar}{12\pi}\tau_{\pm}<T_{\pm\pm}(\phi_{\rm cl})$ (17)
@@
$\left|\braket{T_{IJ}^{2}(x)}_{\rm ren}\right|\leq\epsilon\braket{T_{IJ}(x)}_{\rm ren}^{2}$ (19)
with $\epsilon<1$ for any tetrad. Note that Minkowski with the Poincare invariant vacuum satisfies this at every point.
@@ II Discussion
In the context of 2d black holes, states for which (18) applies include emblematic cases such as the Boulware vacuum, the Hartle-Hawking state, the Unruh state, as well as the in-state (representing the particle creation in the setting of gravitational collapse). This could have been naively anticipated from the Gaussian nature of such states as a general consequence of the Wick theorem. Here we have shown that the expectation remains correct after renormalization. We note that in 4d the a field $\Phi$ in the spherically symmetric modes satisfies the Klein-Gordon equation (for a massive or massless cases) at the black hole horizon if $\Phi=\phi/r$ with $\phi$ satisfying the conformally invariant 2d equation of the field in the previous section. Moreover, the spherical mode contribution to the energy momentum tensor at the black hole horizon, for the relevant component encoding the radiation flux across the horizon, is $T^{\rm sph}_{vv}=T^{\rm 2d}_{vv}/r^{2}$ (see Appendix C). Even when the in and Unruh quantum states would change in the generalization to 4d due to the backscattering by the effective potential that is non trivial outside of the horizon, one expects that the previous identification would be qualitatively sensible for the expectation values and fluctuations of the above quantity. This fact, combined with our discussion of the 4-point function OPE in 4d, indicate that the conclusions should remain valid in any dimension (even when explicit evaluation might be difficult).
@@ III Acknowledgments
We are grateful to B. Kay, B. Juarez Aubry for clarifying discussions in early stages of this work, and to M. Knecht and I. Agullo for helpful exchanges. DS acknowledges support from the Proyect PAPIIT No IG100124, and a sabbatical year Fellowship from UNAM, through a PASPA/DGAPA and the hospitality of the Centre de Physique Théorique in Marseille, and the Departament de Física Quantica i Astrofísica Universitat de Barcelona.
Appendix A Four point function in 4d

### v2

@@ Abstract
We analyze the expectation value of the energy–momentum tensor and its fluctuations in quantum field theory on curved spacetimes $\langle T_{ab}\rangle$. A generally accepeted condition for the conceptual consistency of semiclassical gravity, where $\langle T_{ab}\rangle$ represent the sources of the Einstein equations, is that the fluctuations of the energy momentum tensor remain small compared to its expectation value. We study the renormalization of both the energy–momentum tensor $\langle T_{ab}(x)\rangle_{\rm ren}$ and the fluctuation tensor $\langle T_{ab}(x)T_{cd}(x)\rangle_{\rm ren}$ for suitable Hadamard states, using the operator product expansion for a free scalar field on a fixed curved background. We show that states (usually referred to as ‘squeezed vacua’)—arising naturally in black hole evaporation and in inflationary cosmology—fail to satisfy the natural semiclassicality criterion.
@@
In the absence of a consistent and well-defined theory of quantum gravity, semiclassical gravity—defined by Einstein’s equations coupled to the expectation value of the energy–momentum tensor of matter in a quantum state—is expected to provide useful insights into phenomena at the interface between gravity and quantum physics. Even though it is not entirely clear whether the dynamical equations of semiclassical gravity can be given a fully consistent interpretation—owing to questions about well-posedness in generic situations stemming from their higher-derivative character Flanagan:1996gw (see, Juarez-Aubry:2022qdp for an approach that appears as a promising path to deal with those issues, and including the initial value formulation of the theory) —the semiclassical framework nonetheless plays a central role in conceptual discussions. These range from the physics of black-hole evaporation via Hawking radiation, to analyses of the generalized second law, the fate of information, and various issues related to the emergence of structure in cosmology.
In order to explore the regime of applicability of semiclassical gravity, it is natural to consider the quantum uncertainties of the energy–momentum tensor in comparison to the expectation value of that same quantity. The issue would be straightforward were it not for the fact that both the expectation value and the fluctuations are divergent notions that require suitable renormalization in order to yield finite and meaningful results.
The renormalization of expectation values of operators that are quadratic in the fundamental fields and their derivatives (such as the energy–momentum tensor of free fields) is customarily based on the Hadamard subtraction. An axiomatic analysis proposed by Wald Wald:1995yp assures that the recipe is unique up to the higher order curvature terms which are naturally expected to be suppressed by appropriate powers of the Planck scale.
@@
While our results agree in general lines with these earlier works and in the specific settings they considered, our approach offers a systematic and fully covariant framework that is, in principle, applicable to general spacetimes and general (suitably tamed Hollands:2009bke ) states. Moreover, the OPE perspective ensures automatically that flat spacetime with matter fields in the Minkowski vacuum state is a consistent solution of semiclassical gravity (see (21)).
@@
${{\mathfs{T}}}_{ab^{\prime}}$ $=$ $\nabla_{a}\nabla_{b^{\prime}}-\frac{1}{2}g_{ab^{\prime}}g^{cd^{\prime}}\nabla_{c}\nabla_{d^{\prime}}-\frac{1}{2}g_{ab^{\prime}}m^{2}$ (3)
with $g_{a{b^{\prime}}}$ representing the parallel transport operators along the unique geodesic connecting the points $x$ and $x^{\prime}$ within a common convex neighbourhood, and $H(x,x^{\prime})$—which in equation (2) corresponds to $C(x,x^{\prime})_{\rm uv}$—is the Hadamard bi-distributional kernel, constructed from geometrically defined quantities, which is state-independent and captures the UV behavior of the (Hadamard) states considered physically meaningful.
@@
When studying the fluctuations of the energy momentum tensor in a given quantum state of the scalar field, the strategy will be to start with the 4-point function $\langle\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi\rangle$, from which we can define multi-tensorial quantity
@@
The OPE is very general and allows for the renormalization of the expectation value of arbitrary products of fields and their derivatives at a same point. However, even when the framework ensures the success of the renormalization program, actual calculations can become difficult in concrete practical applications.
@@
Defining
$\braket{0|\phi(x_{1})\phi(x_{2})|0}_{\rm ren}\equiv\braket{0|\phi(x_{1})\phi(x_{2})|0}-H(x_{1},x_{2}),$ (6)
and
$\!\!\!\!\!\!\!\!\!\braket{0|\phi(x_{1})\phi(x_{2})|0}_{\rm ren}\braket{0|\phi(x_{3})\phi(x_{4})|0}_{\rm ren}+{\rm other\ pairings},{}$
we show, in appendix B, that
@@
Note that this corresponds to the OPE of the 4-point function of a free field on a curved background. This follows from performing the multiple Taylor expansion of the renormalized quantities (those with the subindex “ren”) in (8)—for example writing $\braket{\psi|\phi(x)\phi(x+y)|\psi}_{\rm ren}=\braket{\psi|\phi(x)^{2}|\psi}_{\rm ren}+y^{\mu}\braket{\psi|\phi(x)\partial_{\mu}\phi(x)|\psi}_{\rm ren}+\cdots$. As each of the divergent terms in (8) are canonical, the finite term (7) does not depend on the choice of the vacuum state used in the normal ordering. Furthermore, as the various terms exhibit the appropriate degree of divergence, the arrangement of the terms correspond to the OPE and thus allow identification of the desired renormalized expectation value of the composite operator with no additional ambiguities beyond those already present in the renormalization of energy momentum tensor (see Appendix A).
The renormalized square of the energy momentum tensor can now be given a meaning as
@@
Focusing on the trace, (7) implies that for any squeezed state (defined as a Bogoliubov transformation of an arbitrary vacuum state; see Appendix E for more discussion) such that $\braket{\phi^{2}}_{\rm ren}\not=0$, we have
@@ I Explicit calculation in 2d
$\braket{\psi|T_{\pm\pm}(x)|\psi}_{\rm ren}=\braket{\psi|:T_{\pm\pm}:_{x}|\psi}-\frac{\hbar}{12\pi}\tau_{\pm}(\Omega),$ (11)
@@
$\tau_{\pm}(\Omega)\equiv\left(\partial_{\pm}\log(\Omega)\partial_{\pm}\log(\Omega)-\partial^{2}_{\pm}\log(\Omega)\right),$ (12)
@@
where, as in (I), we denote by $_{X}$ the normal ordered prescription based on the vacuum notion constructed out of the positive frequency notion associated with freely falling null coordinates $X^{\pm}$. In Appendix C we explicitly show that the previous procedure is equivalent to the OPE perspective advocated in Hollands:2023txn . We also show that the relevant renormalized fluctuations are given by (see (C.2))
@@
$\frac{\hbar(5+\sqrt{26})}{12\pi}\tau_{\pm}<T_{\pm\pm}(\phi_{\rm cl})$ (18)
when $\tau_{\pm}>0$ and
$\frac{\hbar(5-\sqrt{26})}{12\pi}\tau_{\pm}<T_{\pm\pm}(\phi_{\rm cl})$ (19)
@@
$\left|{\Delta T_{IJ}^{2}(x)}_{\rm ren}\right|\leq\epsilon\braket{T_{IJ}(x)}_{\rm ren}^{2}$ (21)
with $\epsilon<1$ for any tetrad  [^1] One might naively worry about the fact that ambiguities in the renormalization of the quantities involved in the previous discussion might make the criterion (21) ill defined. Such ambiguous contributions appear as higher order modifications of Einstein’s equations of the form $G_{ab}=\ell^{2}_{p}\braket{\tilde{T}_{ab}(x)}_{\rm ren}+\ell^{2}_{p}(\alpha\nabla_{a}\nabla_{b}{R}+\cdots).$ The natural suppression of these terms is ensured by the requirement that both curvatures and their rate of change be small compared with $\ell_{p}$.. Note that Minkowski with the Poincare invariant vacuum satisfies this at every point.
@@ II Discussion
In the context of 2d black holes, states for which (20) applies include emblematic cases such as the Boulware vacuum, the Hartle-Hawking state, the Unruh state, as well as the in-state (representing the particle creation in the setting of gravitational collapse). This could have been naively anticipated from the Gaussian nature of such states as a general consequence of the Wick theorem. Here we have shown that the expectation remains correct after renormalization. We note that in 4d the field $\Phi$ in the spherically symmetric modes satisfies the Klein-Gordon equation (for a massive or massless cases) at the black hole horizon if $\Phi=\phi/r$ with $\phi$ satisfying the conformally invariant 2d equation of the field in the previous section. Moreover, the spherical mode contribution to the energy momentum tensor at the black hole horizon, for the relevant component encoding the radiation flux across the horizon, is $T^{\rm sph}_{vv}=T^{\rm 2d}_{vv}/r^{2}$ (see Appendix D). Even when the in and Unruh quantum states would change in the generalization to 4d due to the backscattering by the effective potential that is non trivial outside of the horizon, one expects that the previous identification would be qualitatively sensible for the expectation values and fluctuations of the above quantity. This fact, combined with our discussion of the 4-point function OPE in 4d, indicate that the conclusions should remain valid in any dimension (even when explicit evaluation might be difficult).
@@ III Acknowledgments
We are grateful to B. Kay, B. Juarez Aubry for clarifying discussions in early stages of this work, and to I. Agullo, L. Freidel, M. Knecht, and E. Verdaguer for helpful exchanges. DS acknowledges support from the Proyect PAPIIT No IG100124, and a sabbatical year Fellowship from UNAM, through a PASPA/DGAPA and the hospitality of the Centre de Physique Théorique in Marseille, and the Departament de Física Quantica i Astrofísica Universitat de Barcelona. This work was made possible through the support of the WOST, WithOut SpaceTime project (https://withoutspacetime.org), supported by Grant ID63683 from the John Templeton Foundation (JTF). The opinions expressed in this work are those of the author(s) and do not necessarily reflect the views of the John Templeton Foundation.
Appendix A The OPE and the uncertainties in the energy momentum tensor
In order to evaluate the uncertainties of the energy momentum tensor we must compute the quantity $\braket{\psi|T_{ab}(x)T_{cd}(x)|\psi}_{\rm ren}$. We approach this task by considering the OPE philosophy and focus on the coincidence limit of the quantity $\braket{\psi|T_{ab}(x_{1})T_{cd}(x_{2})|\psi}_{\rm ren}$. Here we give a brief description of the formalism underlying the procedure which is based on Hollands:2009bke , and applied to the specific case of the energy momentum tensor for a free scalar field in 4 dimensions. The OPE in this case yields
${T}_{a_{1}b_{1}}(x_{1}){T}_{c_{2}d_{2}}(x_{2})=C^{(8)}_{a_{1}b_{1},c_{2}d_{2}}(x_{1},x_{2},y){1}+\sum_{i=1}^{\infty}C^{a^{1}_{y}a^{2}_{y}\cdots a_{y}^{i};(8-i)}_{a_{1}b_{1},c_{2}d_{2}}(x_{1},x_{2},y){O}^{(i)}_{a^{1}_{y}a^{2}_{y}\cdots a_{y}^{i}}(y),$ (22)
where $C^{a^{1}_{y}a^{2}_{y}\cdots a_{y}^{i};(8-2i)}_{a_{1}b_{1},c_{2}d_{2}}(x_{1},x_{2},y)$ are three-tensorial distributions in the tangent planes at $x_{1}$, $x_{2}$, and $y$ with abstract indices $a_{1},b_{1}$, $a_{2},b_{2}$ and $a^{1}_{y}a^{2}_{y}\cdots a_{y}^{i}$ in the three corresponding tangent planes, and $O^{(d)}_{\cdots}$ denotes operators with scaling dimension $d$. For example the supra-index (8i)(8-i) and $(i)$ denote the scaling dimension of the tensor valued distributions and operators in appearing in the OPE, which coincide with the power counting dimension in the present case as we are dealing with free fields (no anomalous dimensions involved). Thus coefficients scale as $1/|\sigma|^{(8-i)}$ where $\sigma$ is the geodesic distance between the points (see Hollands:2009bke for more details). The OPE is meant to be used in evaluating expectation values calculated with the physical states of the theory (see Hollands:2009bke for details) and for points as they approach coincidence. One can write the first few terms with positive scaling dimension, namely
${T}_{a_{1}b_{1}}(x_{1}){T}_{a_{2}b_{2}}(x_{2})=$ (23)
$=C^{(8)}_{a_{1}b_{1},c_{2}d_{2}}(x_{1},x_{2},y){\bf 1}+C^{a_{y}b_{y};(4)}_{a_{1}b_{1},a_{2}b_{2}}(x_{1},x_{2},y){O}^{(4)}_{a_{y}b_{y}}(y)+C^{a_{y}b_{y}c_{y}d_{y};(0)}_{a_{1}b_{1},a_{2}b_{2}}(x_{1},x_{2},y){O}^{(8)}_{a_{y}b_{y}c_{y}d_{y}}(y)+\cdots{}$
[…]
$\braket{\psi|{T}_{a_{1}b_{1}}(x){T}_{a_{2}b_{2}}(x)|\psi}_{\rm ren}\equiv\lim_{x_{1},x_{2}\to x}\,C^{a_{y}b_{y}c_{y}d_{y};(0)}_{a_{1}b_{1},a_{2}b_{2}}(x_{1},x_{2},x)\braket{\psi|{O}^{(8)}_{a_{y}b_{y}c_{y}d_{y}}(x)|\psi}$ (24)
Thus, the OPE provides a recipe for the definition of the quantity of interest, namely the renormalized expectation value, in a suitable physical state of the theory (i.e. the states for which the OPE is valid), of the composite operator $T_{ab}(x)T_{cd}(x)$, just as the Hadamard substraction provides a definition of the renormalized expectation value for a Hadamard state of the composite operator $T_{ab}(x)$.
The axioms of the OPE imply that one can obtain (24) by starting from the OPE corresponding to the product of four scalar field operators. More precisely, as algebraic (and by considering suitable limits also differential) relations are required to be preserved in the corresponding expansion, thus we could have focused on the four point function, instead of looking at the energy momentum tensor directly, and would then have arrived at the desired expression using its corresponding OPE evaluated on the state of interest (as in eq. (8)). Explicitly, starting from
$\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})=C^{(4)}(x_{1},x_{2},x_{3},x_{4};y){\bf\hat{1}}+C^{(2)}(x_{1},x_{2},x_{3},x_{4};y){\phi^{2}}(y)+C^{(0)}(x_{1},x_{2},x_{3},x_{4};y)\phi^{4}(y)+{}$
$+C^{ab;(0)}(x_{1},x_{2},x_{3},x_{4};y)\partial_{a}\phi(y)\partial_{b}\phi(y)+C^{ab;(-2)}(x_{1},x_{2},x_{3},x_{4};y)\phi^{2}(y)\partial_{a}\phi(y)\partial_{b}\phi(y){}$
$+C^{abcd;(-4)}(x_{1},x_{2},x_{3},x_{4};y)\partial_{a}\phi(y)\partial_{b}\phi(y)\partial_{c}\phi(y)\partial_{d}\phi(y)+\cdots,$ (25)
one has that
$\braket{\psi|{T}_{a_{1}b_{2}}(x){T}_{a_{3}b_{4}}(x)|\psi}_{\rm ren}=\lim_{x_{1},x_{2},x_{3},x_{4}\to x}\,\underbrace{\left({\mathfs{T}}_{a_{1}b_{2}}{\mathfs{T}}_{a_{3}b_{4}}\braket{\psi|\phi(x)\phi(x)\phi(x)\phi(x)|\psi}\right)}_{d=8\ \rm term\ in\ the\ OPE\ expansion},{}$
with ${\mathfs{T}}_{ab^{\prime}}$ defined in (3). Note that $C^{(0)}(x_{1},x_{2},x_{3},x_{4};y)$, $C^{ab;(-2)}(x_{1},x_{2},x_{3},x_{4};y)$, and $C^{abcd;(-4)}(x_{1},x_{2},x_{3},x_{4};y)$ contribute to the final result. The general expression in four dimensions can be expected to be rather cumbersome yet available given the four and two point correlation functions in the given states (as shown in equations (7) and (8)). We illustrated in this work the general approach explicitly using the simple case of a massless scalar file in 2 dimensions yet the scheme and the conclusions are general. Indeed, higher powers of the energy momentum tensor can be renormalized with the same framework.
Up to some subtle issues that we will not discuss here, the requirement that the l.h.s. of (23) be divergence-less on both points implies that each one of the coefficients on its own must be divergence free in both points as well, as there is no general possibility of cancellations between the terms due to the independence of the operators $\hat{O}^{(i)}(y)$ occurring in the expansion. That property, in turn, suggests that the ambiguities occurring in the identification of the coefficients are closely related to, and appear not to exceed, those arising in the standard renormalization of $\langle T_{ab}\rangle$ itself.
A formal link between the renormalization ambiguities of the expectation value of the quantities of interest here and the ambiguities in the effective action can be established via the path integral formulation. The expectation value of the energy momentum tensor is given by
$\braket{T_{ab}(x)}\equiv\frac{i\hbar}{Z\sqrt{|g(x)|}}\frac{\delta}{\delta g^{ab}(x)}\underbrace{\left(\int{\mathfs{D}}g_{\mu\nu}{\mathfs{D}}\phi\exp\left(\frac{i}{\hbar}S_{M}[g_{\mu\nu},\phi]\right)\right)}_{\equiv Z}.$ (26)
where $S_{M}[g_{\mu\nu},\phi]$ is the matter action defined as the full action minus the (bare) Einstein-Hilbert action, i. e. , it contains all the terms leading to possible higher curvature corrections to Einstein’s equations (information about the quantum state is meant to be encoded in the boundary conditions of the functional integral). The renormalization ambiguities in the $\braket{T_{ab}(x)}$ mentioned in the introduction (that follow from Wald’s axioms Wald:1995yp ) are thus connected to the renormalization of the parameters in $S_{M}[g_{\mu\nu},\phi]$ like the Newton and cosmological constants as well as the higher curvature couplings like $\alpha\sqrt{|g|}R^{2}+\beta\sqrt{|g|}R_{ab}R^{ab}+\cdots$ appearing in the Lagrangian defining $S_{M}[g_{\mu\nu},\phi]$. Similarly, we can study second variations from which we get
$\frac{-\hbar^{2}}{Z\sqrt{|g(y)|}}\frac{\delta}{\delta g^{cd}(y)}\frac{\delta}{\sqrt{|g(x)|}\delta g^{ab}(x)}\left(\int{\mathfs{D}}g_{\mu\nu}{\mathfs{D}}\phi\exp\left(\frac{i}{\hbar}S_{M}[g_{\mu\nu},\phi]\right)\right)=$ (27)
$=\underbrace{\frac{1}{Z}\left(\int{\mathfs{D}}g_{\mu\nu}{\mathfs{D}}\phi\,T_{ab}(x)T_{cd}(y)\exp\left(\frac{i}{\hbar}S_{M}[g_{\mu\nu},\phi]\right)\right)}_{\equiv\braket{T_{ab}(x)T_{cd}(y)}}+\overbrace{\frac{i\hbar}{Z}\left(\int{\mathfs{D}}g_{\mu\nu}{\mathfs{D}}\phi\,\frac{\delta T_{ab}(x)}{\sqrt{|g(y)|}\delta g^{cd}(y)}\exp\left(\frac{i}{\hbar}S_{M}[g_{\mu\nu},\phi]\right)\right)}^{{\rm contact\ term}=\delta(x,y)\braket{O_{abcd}(y)}+\cdots}.$
This implies that—apart from singular contact terms that would not contribute to $\braket{T_{ab}(x)T_{cd}(x)}_{\rm ren}$—the finite operator expectation value $\braket{T_{ab}(x)T_{cd}(x)}_{\rm ren}$ shares the ambiguities with $\braket{T_{ab}(x)}_{\rm ren}\braket{T_{cd}(x)}_{\rm ren}$, which thus should cancel in the expression of $\Delta(T_{ab}T_{cd})(x)$.
Appendix B Four point function in 4d